OpenStudy (jamanorthi):

Help plzzzz...!!! The membership fee for joining a camping association is $45. A local campground charges members $35 per night and charges non-members $40 per night for a campsite. After how many nights of camping is the total cost for members and non-members, including membership fee, the same as the total cost for non-members???

OpenStudy (texaschic101):

n stands for # of nights 45 + 35n = 45 + 40n I must be doing something wrong because this will not come out right. I set the members fees and non-members fees (both including membership fees) equal to each other. I do not think they will ever be equal.

OpenStudy (jamanorthi):

yeah

OpenStudy (texaschic101):

@TuringTest ...help

OpenStudy (jamanorthi):

the membership fee doesn't go with non members

OpenStudy (jamanorthi):

so its 45+35n=40n

OpenStudy (texaschic101):

yes..if membership fee does not go with non-members

OpenStudy (jamanorthi):

yeah

OpenStudy (jamanorthi):

i got it now. thanks

OpenStudy (texaschic101):

45 + 35n = 40n 45 = 40n - 35n 45 = 5n 9 = n check.. 45 + 35(9) = 40(9) 45 + 315 = 360 360 = 360 (correct) they will cost the same at 9 nights

OpenStudy (texaschic101):

sorry about the mistake earlier...